# EmoDynamiX: Emotional Support Dialogue Strategy Prediction by Modelling MiXed Emotions and Discourse Dynamics

### v1

@@ Appendix E Baselines
Appendix F Hyperparameter Settings
For the pre-training of the ERC module, we set the learning rate to 2e-5, with 500 warm-up steps and a weight decay of 1e-3, training the model for 12,000 steps. For the pre-training of the discourse parser, we adhered to the hyperparameter settings detailed in Chi and Rudnicky (2022). All training procedures were conducted on a single Nvidia Tesla P100 GPU.
Appendix G Impact of the Initialization of $\tau$
@@
Figure 7: Impact of extractive few-shot learning on ChatGPT. Left: predictions of 0 shot; right: predictions of 2 shot.
Figure 8: Impact of extractive few-shot learning on LLaMA-70B. Left: predictions of 0 shot; right: predictions of 2 shot.
@@
Appendix H Preference Bias of LLMs
Further analysis on ESConv indicates that in-context learning does not always positively impact large language models (LLMs) in mitigating their inherent preference bias in dialogue strategy prediction. As illustrated in Figure 7, ChatGPT (0-shot) and ChatGPT (2-shot) demonstrate different outcomes. ChatGPT effectively understands the meanings of strategies by looking at definitions alone. However, when provided with examples, whether randomly selected or extracted using sentence-BERT, its bias was actually reinforced. Conversely, the situation is different with LLaMA3-70B. As shown in Figure 8, LLaMA3-70B relied heavily on a few strategies in the 0-shot setting, indicating its limited ability to understand dialogue strategies from definitions alone. However, its strong in-context learning ability helped mitigate this issue, as it began to predict more diverse dialogue strategies when given extracted examples. It is important to note, however, that randomly selecting examples had a negative effect on LLaMA3-70B, as demonstrated in Table 1.
Appendix I Confusion Matrix of EmoDynamiX on ESConv
See Figure 9.

### v2

@@ Appendix D Baselines
MultiESC (Cheng et al., 2022) is a specialized ESC framework based on BART. It features a look-ahead strategy planning mechanism inspired by A* search algorithm to maximize the expected user feedback.
@@
TransESC (Zhao et al., 2023) is a specialized ESC framework built upon BlenderBot. It models dialogue state transitions using a graph-based approach and integrates emotion recognition as an additional training objective, utilizing ground-truth emotion labels predicted by an off-the-shelf ERC model. TransESC’s modeling of the user’s emotional state also leverages commonsense knowledge from COMET.
Appendix E Implementation of the Preference Bias Score
Preference $p_{i}$ indicates the degree to which the model favors strategy $i$ over others. It is calculated iteratively using the confusion matrix according to the following formula:
$p_{i}^{\prime}=\frac{\sum_{j}(w_{ij}p_{j})/(p_{i}+p_{j})}{\sum_{j}w_{ji}/(p_{i}+p_{j})}$ (12)
Here, $p_{i}^{\prime}$ denotes the updated preference for strategy $i$ in the next iteration, and $w_{ij}$ represents the frequency with which the model predicts strategy $i$ when the actual ground truth is strategy $j$. Initially, all preferences $p_{i}$ are set to 1. In our implementation, we perform 20 iterations of this process.
Preference Bias is the standard deviation of $p$:
$\mathcal{B}=\sqrt{\frac{\sum_{i=1}^{N}(p_{i}-\overline{p})^{2}}{N}}$ (13)
Appendix F Output Statistics of the ERC Module
Table 3 presents the output statistics of our pre-trained ERC module on the ESConv test set, alongside a comparison with the original label distribution of DailyDialog. Notably, the dialogue scenes in ESConv exhibit a higher emotional intensity compared to those in DailyDialog.
ESConv DailyDialog
Anger 1.83 0.99
Disgust 0.70 0.34
Fear 0.61 0.17
Joy 20.17 12.51
Sadness 17.17 1.12
Surprise 0.31 1.77
Neutral 59.22 83.10
Table 3: Comparison between the output label distribution of our ERC module on ESConv and the original label distribution of DailyDialog.
Appendix G Implementation Details for Submodules
Appendix H Hyperparameter Settings
For the pre-training of the ERC module, we set the learning rate to 2e-5, with 500 warm-up steps and a weight decay of 1e-3, training the model for 12,000 steps. For the pre-training of the discourse parser, we adhered to the hyperparameter settings detailed in Chi and Rudnicky (2022). All training procedures were conducted on a single Nvidia GeForce RTX 4090 GPU.
@@ Appendix H Hyperparameter Settings
Appendix I Impact of the Initialization of $\tau$
@@
Appendix J Confusion Matrix of EmoDynamiX on ESConv
See Figure 8.